### v2

# The Coase Theorem and Ideal Exchanges

Abstract.
This paper offers a proof of the Coase theorem by formalizing the notion of ideal exchanges.
Key words and phrases:
Coase’s Theorem, microeconomics, externalities, Pareto optimality, invariance, convergence, Ronald Coase, Philosophy of Economics.
2020 Mathematics Subject Classification:
Primary 91B02, 03E75, 03A10, 91-03.
1. Introduction
An externality can be defined as an effect, often negative, that is imposed on a third-party through a private activity. An instance of this problem occurs when non-smokers and smokers are within each other’s vicinity; the non-smoker is adversely affected by the smoker’s consumption of cigarettes through no fault of their own. The economic and philosophical significance of the existence of externalities lies in the challenge it poses to the classical position that a free market economy is capable of allocating resources in a socially efficient manner; if the social costs of private consumption are not reflected in the market price of the product in question, then consumers would be incentivized to consume the good at a level that is beyond socially optimal. A comprehensive overview of the economic explanations surrounding externalities and their historical origins can be found in [3].
Prior to Ronald Coase’s work, The Problem of Social Cost, the conventional view—which Coase himself attributes to Arthur Pigou [1, p. 1]—held that externalities could be corrected through state intervention: namely, through taxes or subsidies that are designed to deter people from either over-consuming goods associated with negative externalities or under-consuming goods associated with positive externalities [2]. Coase reoriented the debate from one that relied solely on governmental regulation to one that recognized the potential efficiency of private negotiation in resolving externalities. In particular, Coase reasoned that externalities resulted from ill-defined property rights; provided that property rights are well-defined and each party could negotiate with the other without hindrance, resources would be distributed away from those who value them less and towards those who value them more irrespective of how the rights are initially distributed [1, p. 8, para. 2]. This proposition, although informal, was characterized as a theorem in George Stigler’s Theory of Price:
The Coase theorem thus asserts that under perfect competition private and social costs will be equal. It is a more remarkable proposition to us older economists who have believed the opposite for a generation, than it will appear to the younger reader, who was never wrong, here. [4]
Returning to the initial example, if the non-smoker has the right to clean air, the smoker would have to compensate them to smoke in shared spaces. Conversely, if the smoker has the right to smoke, the non-smoker might compensate the smoker to refrain from smoking near them. Hence, regardless of which party initially holds the right, the outcome of these negotiations will be efficient, as the final allocation of rights will reflect the true costs and benefits to both sides.
Stigler’s recognition of the potential mathematical standing of what is now widely termed the Coase theorem invites us to examine the economic proposition precisely, though an exact formulation is not present in Coase’s original work. The prevailing interpretation of the theorem adopted by this paper can be stated as follows: if property rights are well-defined and there are no transaction costs, then rational agents would arrive at an optimal distribution of resources on their own accord. We find this interpretation in Stigler [4], Nutter [5], Demsetz [6], and Coase himself [1]. Yet, again, we are confronted with the difficulty of constructing a precise interpretation of informal notions such as the property of being “well-defined” or constituting a “transaction cost”. Additionally, we may interpret the notion of “arriving at an optimal outcome” as a declaration of the existence of some ultimate, optimal distribution such that it is a necessary outcome of every possible initial distribution; alternatively, we may consider that there are multiple ultimate distributions for each initial distribution that fall within the set of outcomes characterized as being optimal.
This paper endeavors to formalize interpretations of the Coase theorem as set-theoretic propositions. We offer the philosophical motivations for doing this in Section 2 and construct the formal aspects of the theorem thereafter. Where necessary, explanatory remarks are offered to justify the formalization. We show that certain interpretations of the theorem are false whereas one is true if we admit the notion of ideal exchanges. Lastly, we conclude by offering some philosophical remarks on these results in Section 6.
2. Preliminaries
Contemporary methods of illustrating the Coase theorem are either informal or rely on mathematical structures that are unnatural within the discipline of economics. We shall concern ourselves with the latter. One such illustration can be obtained by employing the standard techniques of an indifference curve analysis, which are readily available in any microeconomics textbook.
Consider the existence of two types of goods, $x$ and $y$, such that $Q_{x}$ denotes the quantity of the former and $Q_{y}$ denotes the quantity of the latter. It is standard to assume that each 2-tuple $(Q_{x},Q_{y})$ is present in $\mathbb{R}^{+}\times\mathbb{R}^{+}$ or, equivalently, +$\mathbb{R}^{2}_{+}$. [^1] We include 00 in $\mathbb{R}^{+}$. We proceed by declaring the existence of a utility function for some agent $A$ such that $\mathcal{U}_{A}:\mathbb{R}^{2}_{+}\to\mathbb{R}^{+}$ to model the notion that it is possible to map pairs of quantities of goods to some level of utility. [^2] It would seem that happiness is a crude synonym for utility, since the latter is often portrayed as a quantifiable notion. An agent is indifferent between two outcomes iff they render the same level of utility. In other words, an indifference curve is a set of all the 2-tuples $(Q_{x},Q_{y})\in\mathbb{R}^{2}_{+}$ such that $\mathcal{U}_{A}(Q_{x},Q_{y})=c$, where $c$ is a constant in $\mathbb{R}^{+}$. Furthermore, it is standard for an economist to assume that each indifference curve carries important characteristics. Intuitively speaking, it would be peculiar for an indifference curve to contain $(Q_{x},Q_{y})$ and $(Q_{x}^{\prime},Q_{y}^{\prime})$ if $Q_{x}<Q_{x}^{\prime}$ and $Q_{y}<Q_{y}^{\prime}$, since that would imply that the utility of the agent remains constant despite an increase in the presence of both goods. It follows, therefore, that all indifference curves must be monotonic and decreasing. A more refined assumption would be that indifference curves should model the empirical observation of diminishing marginal utility, that is, every additional unit of utility gained from each additional unit of the good decreases, but never goes below zero. [^3] If it did, then that would imply that an additional unit of the good resulted in a loss of utility, which is not a standard assumption in microeconomic theory. Thus, indifference curves are often portrayed as being convex to the agent’s origin.
Let us now consider an exchange economy where there are two goods ($x$ and $y$) and two agents ($A$ and $B$). We represent this with the set $\mathbb{E}=[0,Q_{x}]\times[0,Q_{y}]$, where $\mathbb{E}$ stands for an Edgeworth Box. Every point in the box effectively forms a partition of the goods among both agents. For example, the point $(0,0)_{B}=(Q_{x},Q_{y})$ in Figure 1 is a distribution of resources where $A$ is in possession of everything and $B$ is in possession of nothing; the opposite is true at $(0,0)_{A}=(0,0)$.
Figure 1: An Edgeworth Box $(0,0)_{A}$$x_{A}$$y_{A}$$(0,0)_{B}$$x_{B}$$y_{B}$$\omega_{A}$$\omega_{B}$$\omega_{A}^{\prime}$$\omega_{B}^{\prime}$$\Omega$
Consider a family of sets for each agent:
  1. (F1)
    Family of Indifference Curves:
    $\forall c\in\mathbb{R}^{+}\,\Big{[}\big{\{}(x_{A},y_{A})\,|\,\mathcal{U}_{A}(x_{A},y_{A})=c\big{\}}\in\mathcal{F}_{A}\,\land\,\big{\{}(Q_{x}-x_{A},Q_{y}-y_{A})\,|\,\mathcal{U}_{B}(Q_{x}-x_{A},Q_{y}-y_{A})=c\big{\}}\in\mathcal{F}_{B}\Big{]}.$
Three members of $\mathcal{F}_{A}$ appear in Figure 1 as blue curves, while three members of $\mathcal{F}_{B}$ are shown as downward-sloping red curves. A standard definition of the notion of optimality thus follows: if it is not possible to make one agent better off without there being another agent who is made worse off, then the current distribution is Pareto optimal. i.e.,
  1. (P1)
    Pareto Optimality of $(x_{A},y_{A})$:
    $\forall x_{A}^{\prime},y_{A}^{\prime}\text{ s.t. }x_{A}^{\prime}\neq x_{A},\,y_{A}^{\prime}\neq y_{A},\,\frac{\mathcal{U}_{A}(x_{A}^{\prime},y_{A}^{\prime})-\mathcal{U}_{A}(x_{A},y_{A})}{\mathcal{U}_{B}(Q_{x}-x_{A}^{\prime},Q_{y}-y_{A}^{\prime})-\mathcal{U}_{B}(Q_{x}-x_{A},Q_{y}-y_{A})}\in\mathbb{R}_{<0}$
Every Pareto optimal distribution occurs when the indifference curves for both agents are tangent to each other. We can see this by considering the intersection at $\omega_{A}$ or, equivalently, $\omega_{B}$. Since $A$ has an excess of good $x$ and $B$ has an excess of good $y$, it would be possible for $A$ to give up just enough of $x$ and receive just enough of $y$ so that she is indifferent between $\omega_{A}$ and $\Omega$, while $B$’s indifference curve progresses to a point where it is no longer possible to gain any further utility without damaging $A$’s material interests; the opposite is true at $\omega_{A}^{\prime}$. If we assume that exchanges can occur iff they are mutually beneficial, then the negotiated outcome is strictly inside the lens formed by the two indifference curves. Through continual negotiations, the size of the lens shrinks until it is no longer possible to make a mutually beneficial trade. [^4] This is a standard description of how negotiations occur. It is notably present in Buchanan and Tullock’s foundational work on public choice theory, The Calculus of Consent: Logical Foundations of Constitutional Democracy [8, p. 100]. Hence, the ultimate outcome, although different from $\Omega$, would still be Pareto optimal. A stronger variant of the theorem in narrow cases where a medium of exchange is involved was proven by Leonid Hurwicz, who demonstrated that the invariance of the final distribution of a particular resource necessarily depended on the assumption of quasi-linear preferences (i.e., $\mathcal{U}_{A}(Q_{x},m)=v(Q_{x})+m$, where $m$ is a “numeraire” good) [9].
It is a striking fact that the preceding methods are standard in microeconomic theory yet are obviously mathematically unnatural for the purposes of modeling economic behavior. Francis Ysidro Edgeworth, after whom the Edgeworth box is named, introduced the concept of indifference curves in Mathematical Psychics (MP), a foundational work in contemporary microeconomic theory [10, p. 28]. The broad purpose of MP was to apply mathematical techniques to the moral sciences; in doing so, Edgeworth introduced utility measurements by assigning real numbers to represent individuals’ levels of satisfaction or happiness, facilitating the comparison and aggregation of utility across different people. Furthermore, MP readily and implicitly utilizes the concept of a continuum—suggesting infinitely divisible goods or utility levels—without being concerned with foundational aspects of real analysis. On a historical note, Dedekind’s Stetigkeit und irrationale Zahlen was published in 1872, merely 9 years before MP [11]; it is therefore likely that Edgeworth was unaware of these developments in the philosophical foundations of mathematics and their effects on the mathematical foundations of economics.
A mathematically bizarre example of such an effect is as follows; suppose that the utility any agent derives from the consumption of a unit of good $x$ is precisely $\sqrt{2}$. For the purposes of exchanging the good, agent $A$ would like to utilize a “Dedekind cutter” to mint a medium of exchange such that the numerical value of the medium is precisely identical to that of the utility derivable from the consumption of the good. Although the process by which coins are minted is initially costless, the cost becomes prohibitively high the moment the agent endeavors to mint a sum of coins whose collective face value, when squared, exceeds $2$; moreover, each coin can only express positive rational numbers. Suppose that we have another agent, $B$, who is in possession of the good. Evidently, both agents are willing to perform an exchange, yet they are unable to do so, because one cannot express an irrational number in terms of a finite sum of rational ones. We now consider another mathematically bizarre exchange economy consisting of two agents and two types of goods where the utility anyone can derive from some quantity of good $x$ is a rational multiple of $\sqrt{2}$ and that derivable from some quantity of $y$ is a rational multiple of $\sqrt{\pi}$. Suppose $A$ were in possession of all $x$ and $B$ were in possession of all $y$ and both agents are equally unwilling to allow the other to get more than what is exchanged in return. If $B$ tried to obtain one unit of $x$, she would have to divide her good to an infinite extent, namely, $\sqrt{\frac{2}{\pi}}$, in order for an exchange to occur. It would seem that such exchange economies would paralyze in the absence of “Dedekind cutters”, which themselves are discretely exchangeable; without a sufficiently rigorous description of how bargaining occurs, optimal allocations, despite being guaranteed to exist, are not guaranteed to be reached. Finally, consider an economy where two agents are present and where the first is in possession of an infinitely divisible good, literally the continuum represented by the set $[0,1]$, and the second has the continuum $[2,3]$; it would seem that resources are scarce because only two continua are present. But if we take one real number from $[0,1]$ and give it to the second agent, both agents are indifferent. Agent 1 is not made worse off since $\mathfrak{c}-1=\mathfrak{c}$. Likewise, $\mathfrak{c}+1=\mathfrak{c}$, so the distribution is not Pareto optimal, and permanently so, since we can always make one agent better off without making another worse off.
With this said, it would be natural ask whether it is even possible to axiomatize the foundations of economic theory in terms of logic or mathematics. Should an economic axiom be intrinsically accepted on the basis that it is an empirical observation, such as the assumption of self-interest, diminishing marginal returns or monotonic preferences, despite the fact that it produces unobservable or unfalsifiable conclusions? Or should it be extrinsically accepted because it is capable of accounting for some definite, observable phenomena, despite being intrinsically false? At an academic conference on the discovery of forcing, Paul Cohen noted that
The attempts to formalize mathematics and make precise what the axioms are were never thought of as attempts to explain the rules of logic, but rather, to write down these rules and axioms which appeared to correspond to what contemporary mathematicians were using. [12, p. 1074]
We adopt a similar position with regards to this endeavor to axiomatize a narrow scion of the foundations of economic theory. We certainly do not claim that we are uncovering fundamental, irrefutable truths about economic behavior; we endeavor only to capture and make precise a particular model of the axioms (or hypotheses) that, if true, would produce a set of interesting and non-trivial results; the Coase theorem is one such result.
3. Foundations
Let $\mathbb{A}$ be a set of agents of cardinality $n\in\mathbb{N}$ and $R$ be a set of resources of cardinality $m\in\mathbb{N}$. The distribution of resources across each instance of trade $t\in\mathbb{N}$ can be characterized as a mapping $\mathcal{I}_{t}:\mathbb{A}\to\mathcal{P}(R)$ such that the following two conditions are satisfied:
  1. (A1)
    Well-defined Ownership Rights:
    $\forall\,t\in\mathbb{N}\,\forall\,i\neq j\in[0,n-1]\,\big{(}\mathcal{I}_{t}(a_{i})\cap\mathcal{I}_{t}(a_{j})=\varnothing\big{)}.$
  2. (A2)
    Absence of Unclaimed Resources:
    $\forall\,t\in\mathbb{N}\,\bigg{[}\bigcup_{i=0}^{n-1}\mathcal{I}_{t}(a_{i})=R\bigg{]}.$
Remark 3.1.
These assumptions allow $R$ to be partitioned among $n$ agents. Furthermore, they effectively ensure that no transaction costs are present, since all resources are always being possessed by some agent.
The preferences of each agent can be represented as a mapping $W:\mathbb{A}\to\mathcal{P}\big{(}\mathcal{P}(R)\times\mathcal{P}(R)\big{)}$ such that they are characteristically rational and therefore satisfy the following conditions:
  1. (B1)
    Asymmetry:
    $\forall\,i\in[0,n-1]\,\forall\,A,B\in\mathcal{P}(R)\,\big{[}(A,B)\in W(a_{i})\implies(B,A)\not\in W(a_{i})\big{]}.$
  2. (B2)
    Transitivity:
    $\forall\,i\in[0,n-1]\,\forall\,A,B,C\in\mathcal{P}(R)\,\big{[}(A,B)\in W(a_{i})\,\land\,(B,C)\in W(a_{i})\implies(A,C)\in W(a_{i})\big{]}.$
  3. (B3)
    Completeness:
    $\forall\,i\in[0,n-1]\,\forall\,A,B\in\mathcal{P}(R)\,\big{[}(A,B)\in W(a_{i})\,\lor\,(B,A)\in W(a_{i})\big{]}.$
Remark 3.2.
The asymmetry of one’s preferences rests on the position that a rational agent cannot be indifferent between materially distinct outcomes. Transitivity and completeness are standard assumptions in economic theory; the latter ensures that the agent is capable of making meaningful comparisons across all possible options and the former ensures that it is possible to make meaningful inferences about one’s preferences. If completeness were rejected, then there could exist some bundle $A^{\prime}$ such that $a_{q}$ is silent on its existence. This is not rational since it does not prescribe a course of action if $A^{\prime}$ were offered to the agent. If transitivity were rejected, then it is possible for an agent to prefer $B$ over $A$ and $C$ over $B$ without preferring $C$ over $A$. Suppose an offer were made to such an agent to exchange their $A$ for $C$. The offer would be rejected on the grounds that the agent does not have an explicit preference for $C$ over $A$. But the agent admits that $C$ is a superior material outcome relative to $A$ by accepting an exchange with $B$ and then with $C$. It follows that transitivity is an essential and defining quality of rational conduct. [^5] This style of reasoning is similar to a dutch book argument [13].
We now consider two inductive features of $\mathcal{I}_{t}$ that hold for all $t\in\mathbb{N}$:
  1. (C1)
    Double Coincidence of Wants:
    $\exists i\neq j\in[0,n-1]\exists R_{1},R_{2}\in\mathcal{P}(R)\Big{[}\big{[}\big{(}(R_{1},R_{2})\in W(a_{i})\land(R_{2},R_{1})\in W(a_{j})\big{)}\land\big{(}R_{1}\subseteq\mathcal{I}_{t}(a_{i})\land R_{2}\subseteq\mathcal{I}_{t}(a_{j})\big{)}\big{]}$
    $\land\big{[}\big{(}\mathcal{I}_{t}(a_{i}),\big{(}\mathcal{I}_{t}(a_{i})\setminus R_{1}\big{)}\cup R_{2}\big{)}\in W(a_{i})\land\big{(}\mathcal{I}_{t}(a_{j}),\big{(}\mathcal{I}_{t}(a_{j})\setminus R_{2}\big{)}\cup R_{1}\big{)}\in W(a_{j})\big{]}\Big{]}$
    $\implies\exists\,i\neq j\in[0,n-1]\,\exists R_{1},R_{2}\in\mathcal{P}(R)\Big{(}\mathcal{I}_{t+1}(a_{i})=\big{(}\mathcal{I}_{t}(a_{i})\setminus R_{1}\big{)}\cup R_{2}\,\land\,\mathcal{I}_{t+1}(a_{j})=\big{(}\mathcal{I}_{t}(a_{j})\setminus R_{2}\big{)}\cup R_{1}\Big{)}.$
  2. (C2)
    Stagnate in the Absence of Mutually Beneficial Trades:
    $\forall i\neq j\in[0,n-1]\forall R_{1},R_{2}\in\mathcal{P}(R)\Big{[}\big{[}\big{(}(R_{1},R_{2})\not\in W(a_{i})\lor(R_{2},R_{1})\not\in W(a_{j})\big{)}\lor$
    $\big{(}R_{1}\not\subseteq\mathcal{I}_{t}(a_{i})\lor\,R_{2}\not\subseteq\mathcal{I}_{t}(a_{j})\big{)}\big{]}\lor\big{[}\big{(}\mathcal{I}_{t}(a_{i}),\big{(}\mathcal{I}_{t}(a_{i})\setminus R_{1}\big{)}\cup R_{2}\big{)}\not\in W(a_{i})\lor$
    $\big{(}\mathcal{I}_{t}(a_{j}),\big{(}\mathcal{I}_{t}(a_{j})\setminus R_{2}\big{)}\cup R_{1}\big{)}\not\in W(a_{j})\big{]}\Big{]}\implies\forall i\in[0,n-1]\big{(}\mathcal{I}_{t+1}(a_{i})=\mathcal{I}_{t}(a_{i})\big{)}.$
Remark 3.3.
These constructions also preclude transaction costs because no element is lost through the friction of trade.
Lastly, we construct a definition of Pareto optimality.
Definition 3.4 (Distribution).
A distribution of resources is an $n$-tuple of the form $\big{(}\mathcal{I}_{t}(a_{i})\big{)}_{i=0}^{n-1}$ in $\mathcal{P}(R)^{n}$ such that it satisfies all preceding conditions.
Definition 3.5 (Projection Function).
If $\mathfrak{K}$ is a distribution where $\mathfrak{K}=\big{(}\mathcal{I}^{\prime}_{t}(a_{i})\big{)}_{i=0}^{n-1}$, then there exists a function $\hat{\mathfrak{K}}:\mathbb{A}\to\mathcal{P}(R)$ such that $\forall a_{q}\in\mathbb{A}[\hat{\mathfrak{K}}(a_{q})=\alpha\,\Longleftrightarrow\,\mathcal{I}^{\prime}_{t}(a_{q})=\alpha]$.
Definition 3.6 (Strict Distributional Preference).
Given an agent $a_{q}$ and the distributions $\mathfrak{I}=\big{(}\mathcal{I}_{t}(a_{i})\big{)}_{i=0}^{n-1}$, $\mathfrak{K}=\big{(}\mathcal{I^{\prime}}_{t}(a_{i})\big{)}_{i=0}^{n-1}$, we say that the agent strictly prefers $\mathfrak{K}$ over $\mathfrak{I}$ (i.e., ^$\hat{\mathfrak{I}}(a_{q})\prec\hat{\mathfrak{K}}(a_{q})$) iff $\big{(}\mathcal{I}_{t}(a_{q}),\mathcal{I^{\prime}}_{t}(a_{q})\big{)}\in W(a_{q}).$
Definition 3.7 (Pareto Optimality).
Consider the distribution $\mathfrak{I}$. We say that $\mathfrak{I}$ is Pareto optimal iff for all alternative distributions $\mathfrak{K}$, if there exists an agent who strictly prefers $\mathfrak{K}$ over $\mathfrak{I}$, then there must be some other agent for whom $\mathfrak{I}$ is strictly preferred over $\mathfrak{K}$, i.e.,
$\forall\,\mathfrak{K}\neq\mathfrak{I}\in\mathcal{P}(R)^{n}\,\bigg{[}\,\exists\,a_{q}\in\mathbb{A}\,\big{[}\hat{\mathfrak{I}}(a_{q})\prec\hat{\mathfrak{K}}(a_{q})\big{]}\implies\exists\,a_{z}\neq a_{q}\in\mathbb{A}\,\big{[}\hat{\mathfrak{K}}(a_{z})\prec\hat{\mathfrak{I}}(a_{z})\big{]}\bigg{]}.$
4. Secondary Results
Lemma 4.1 (Convergence Lemma).
$\forall\,\big{(}\mathcal{I}_{0}(a_{i})\big{)}_{i=0}^{n-1}\in\mathcal{P}(R)^{n}\,\exists\,t\in\mathbb{N}\,\forall\,k\in\mathbb{N}\Big{[}\big{(}\mathcal{I}_{t}(a_{i})\big{)}_{i=0}^{n-1}=\big{(}\mathcal{I}_{t+k}(a_{i})\big{)}_{i=0}^{n-1}\Big{]}.$
Proof.
Suppose, to the contrary, that there exists some initial distribution $\big{(}\mathcal{I}_{0}(a_{i})\big{)}_{i=0}^{n-1}\in\mathcal{P}(R)^{n}$ such that for all $t\in\mathbb{N}$, there is a $k\in\mathbb{N}$ where $\big{(}\mathcal{I}_{t}(a_{i})\big{)}_{i=0}^{n-1}\neq\big{(}\mathcal{I}_{t+k}(a_{i})\big{)}_{i=0}^{n-1}$. In other words, every “stable” distribution that follows from this initial distribution is temporary. There cannot be a case where a distribution is constant for at least two instances before undergoing a non-trivial change; suppose, to the contrary, that such a phenomenon occurred, then the first distribution at $t$ is either one where no agent is willing to trade with another agent, or where no agent is able to trade with another agent. If a break in stagnation occurs at $t+k$, then both conditions must be satisfied at $t+k-1$, but this is contradictory because the distribution at $t+k-1$ descended from $t$. It follows that $\forall t\in\mathbb{N}\Big{[}\big{(}\mathcal{I}_{t}(a_{i})\big{)}_{i=0}^{n-1}\neq\big{(}\mathcal{I}_{t+1}(a_{i})\big{)}_{i=0}^{n-1}\Big{]}$. By (B1), no exchange can be reversed. Therefore, for indefinitely distinct trades, either the bundles are always different or the agents are always different. Since there are only finitely many resources ($m$) and finitely many agents ($n$), there are finitely many possible distributions ($n^{m}$) that dually comply with (A1) and (A2); hence, the existence of at least one cycle is guaranteed when $t=n^{m}$. [^6] This is due to the pigeonhole principle. By (B2), every cycle is reducible to a reversal of one’s preferences, thereby contradicting asymmetry. It follows that such an initial distribution cannot exist. ∎
Remark 4.2.
It would seem that even in the presence of transaction costs, convergence is guaranteed due to the asymmetric nature of the preferences of finite agents over finite resources.
Proposition 4.3 (Invariance of the Ultimate Outcome).
$\forall\,\big{(}\mathcal{I}_{0}(a_{i})\big{)}_{i=0}^{n-1}\in\mathcal{P}(R)^{n}\,\exists\,t\in\mathbb{N}\,\forall\,k\in\mathbb{N}\,\exists\,\big{(}\alpha_{i}\big{)}_{i=0}^{n-1}\in\mathcal{P}(R)^{n}\Big{[}\big{(}\mathcal{I}_{t+k}(a_{i})\big{)}_{i=0}^{n-1}=\big{(}\alpha_{i}\big{)}_{i=0}^{n-1}\Big{]}.$
Theorem 4.4.
Proposition 4.3 is false.
Proof.
Having previously established Lemma 4.1, it suffices to construct some pair of distinct initial allocations in $\mathcal{P}(R)^{n}$ such that they converge to distinct final allocations. Consider the set of initial allocations in $\mathcal{P}\big{(}\{x\}\big{)}^{n}$ such that $\forall i\in[0,n-1]\Big{(}W(a_{i})=\Big{\{}\big{(}\{\},\{x\}\big{)}\Big{\}}\Big{)}$. Consider an initial allocation where $\exists i\in[0,n-1]\big{(}\mathcal{I}_{0}(a_{i})=\{x\}\big{)}$ and $\forall j\neq i\in[0,n-1]\big{(}\mathcal{I}_{0}(a_{j})=\{\}\big{)}$. By (C2) and 4.1, the distribution is immediately stable and permanent. Now, consider a different initial allocation where $\exists k\in[0,n-1]\big{(}\mathcal{I}_{0}(a_{k})=\{x\}\big{)}$, $k\neq i$, and $\forall j\neq k\in[0,n-1]\big{(}\mathcal{I}_{0}(a_{j})=\{\}\big{)}$. This distribution is also immediately stable and permanent yet it is different from the ultimate distribution where, in lieu of $k$, agent $i$ was in possession of resource $x$. In other words, there exists a pair of cases where the initial distribution produces a non-trivial effect on the ultimate distribution. ∎
Proposition 4.5.
Every ultimate distribution is Pareto optimal.
Theorem 4.6.
Proposition 4.5 is false.
Proof.
Firstly, observe that (B1), (B2) and (B3) characterize a linear ordering of all elements in $\mathcal{P}(R)$. For each individual agent $a_{q}$, we reduce these preferences to a $2^{m}$-tuple through $\mathfrak{P}:\mathbb{A}\to\mathcal{P}(R)^{|\mathcal{P}(R)|}$, where $\mathfrak{P}(a_{q})=\big{(}\alpha_{i}\big{)}_{i=0}^{2^{m}-1}$, such that the following holds:
$\forall a_{q}\in\mathbb{A}\bigg{[}\exists\,(A,B)\in W(a_{q})\Longleftrightarrow\exists\,i\neq j\in[0,2^{m}-1]\big{[}(\alpha_{i}=A\,\land\,\alpha_{j}=B)\,\land\,(i<j)\big{]}\bigg{]}.$
We may now construct an explicit counterexample to 4.5. Suppose that $\mathbb{A}=\{a_{0},a_{1},a_{2}\}$, $R=\{x,y,z\}$, and the following statements are satisfied:
  1. (1)
    $\mathfrak{P}(a_{0})=\big{(}\{\},\{z\},\{x\},\{y\},\{z,x\},\{z,y\},\{x,y\},\{x,y,z\}\big{)}$.
  2. (2)
    $\mathfrak{P}(a_{1})=\big{(}\{\},\{y\},\{z\},\{x\},\{y,z\},\{y,x\},\{z,x\},\{x,y,z\}\big{)}$.
  3. (3)
    $\mathfrak{P}(a_{2})=\big{(}\{\},\{x\},\{y\},\{z\},\{x,y\},\{x,z\},\{y,z\},\{x,y,z\}\big{)}$.
  4. (4)
    $\big{(}\mathcal{I}_{0}(a_{0})=\{x\}\,\land\,\mathcal{I}_{0}(a_{1})=\{z\}\big{)}\,\land\,\mathcal{I}_{0}(a_{2})=\{y\}.$
  5. (5)
    $\mathfrak{K}=\big{(}\mathcal{I^{\prime}}_{t}(a_{i})\big{)}_{i=0}^{2}.$
Notice that no agent has an incentive to trade. Agent $a_{0}$ would be willing to exchange her $x$ for $y$, but $a_{2}$ is not willing to trade her $y$ for $x$; instead, she would prefer to exchange it for $z$. Similarly, $a_{1}$ is unwilling to exchange her $z$ for $y$ and would rather trade it for $x$. However, $a_{0}$ is also unwilling to exchange her $x$ for $z$. By (C2) and 4.1, it follows that this distribution is ultimate. Consider an alternative distribution $\mathfrak{K}$ where $\mathfrak{K}=\big{(}\{y\},\{x\},\{z\}\big{)}$. Since $a_{0}$ now has $y$ as opposed to $x$, she strictly prefers $\mathfrak{K}$ over the initial distribution. This is also true for $a_{1}$ and $a_{2}$. It follows that since no agent is made worse off in this alternative distribution $\mathfrak{K}$, the initial distribution is not Pareto optimal. ∎
5. Primary Result
Suppose we accept every assumption in Section 3 aside from (C1) and (C2). Instead, we shall adopt the following in their place for all $t\in\mathbb{N}$:
  1. (D1)
    Ideal Exchanges: Suppose that $\forall t\in\mathbb{N}\Big{[}\mathfrak{I}_{t}=\big{(}\mathcal{I}_{t}(a_{i})\big{)}_{i=0}^{n-1}\Big{]}.$ An ideal exchange is a situation where
    $\exists\,\mathfrak{K}\neq\mathfrak{I}_{t}\in\mathcal{P}(R)^{n}\,\bigg{[}\,\exists\,a_{q}\in\mathbb{A}\,\big{[}\hat{\mathfrak{I}}_{t}(a_{q})\prec\hat{\mathfrak{K}}(a_{q})\big{]}\,\land\,\forall\,a_{z}\neq a_{q}\in\mathbb{A}\,\big{[}\hat{\mathfrak{K}}(a_{z})\not\prec\hat{\mathfrak{I}}_{t}(a_{z})\big{]}\bigg{]}\implies\big{(}\mathfrak{I}_{t+1}=\mathfrak{K}\big{)}.$
  2. (D2)
    Stagnate if Pareto Optimality is Reached:
    $\forall\,\mathfrak{K}\neq\mathfrak{I}_{t}\in\mathcal{P}(R)^{n}\,\bigg{[}\,\exists\,a_{q}\in\mathbb{A}\,\big{[}\hat{\mathfrak{I}}_{t}(a_{q})\prec\hat{\mathfrak{K}}(a_{q})\big{]}\implies\exists\,a_{z}\neq a_{q}\in\mathbb{A}\,\big{[}\hat{\mathfrak{K}}(a_{z})\prec\hat{\mathfrak{I}}_{t}(a_{z})\big{]}\bigg{]}\implies\big{(}\mathfrak{I}_{t+1}=\mathfrak{I}_{t}\big{)}.$
We may now prove the Coase theorem.
Theorem 5.1 (The Coase Theorem).
Every initial distribution of resources converges to some ultimate distribution that is Pareto optimal, i.e.,
$\forall\,\mathfrak{I}_{0}\in\mathcal{P}(R)^{n}\,\exists\,t\in\mathbb{N}\,\forall\,k\in\mathbb{N}\Bigg{[}\big{[}\mathfrak{I}_{t}=\mathfrak{I}_{t+k}\big{]}\,\land\,\forall\,\mathfrak{K}\neq\mathfrak{I}_{t+k}\in\mathcal{P}(R)^{n}\,$
$\bigg{[}\,\exists\,a_{q}\in\mathbb{A}\,\big{[}\hat{\mathfrak{I}}_{t+k}(a_{q})\prec\hat{\mathfrak{K}}(a_{q})\big{]}\implies\exists\,a_{z}\neq a_{q}\in\mathbb{A}\,\big{[}\hat{\mathfrak{K}}(a_{z})\prec\hat{\mathfrak{I}}_{t+k}(a_{z})\big{]}\bigg{]}\Bigg{]}.$
Proof.
Suppose, to the contrary, that there exists an initial distribution $\mathfrak{I}_{0}$ that either fails to converge to some ultimate distribution or fails to converge to an ultimate distribution that is Pareto optimal. There are two cases: either $\mathfrak{I}_{0}$ is Pareto optimal, or it is not. If $\mathfrak{I}_{0}$ is Pareto optimal, then so is $\mathfrak{I}_{1}$; by (D2), every subsequent distribution must be Pareto optimal, since they are all identical to $\mathfrak{I}_{0}$. On the other hand, if $\mathfrak{I}_{0}$ is not Pareto optimal, then there must exist an alternative distribution $\mathfrak{K}$ for which there is an agent $a_{q}$ who strictly prefers $\mathfrak{K}$ over $\mathfrak{I}_{0}$ without any other agent strictly preferring $\mathfrak{I}_{0}$ over $\mathfrak{K}$. There cannot be a situation where every subsequent distribution remains suboptimal. Suppose, to the contrary, that every subsequent distribution is suboptimal; then for every subsequent distribution, it is always possible to make someone better off without there being someone who is worse off, thereby contradicting the assumption of finite resources. It follows that the distribution must converge, and by (D2), only does so when the outcome is Pareto optimal. ∎
6. Philosophical & Concluding Remarks
Firstly, observe that Propositions 4.3 and 4.5 are formal variants of how the Coase theorem can be informally interpreted. The contention that every initial distribution converges to one unique outcome is a particularly strong claim; it is generally false because if every person endeavors to maximize what they own, and for each initial distribution there is one distinct agent who is in possession of all resources, then the ultimate outcome is both Pareto optimal and halts on this initial input irrespective of what this input may be. The invariance of the ultimate distribution is a guaranteed phenomenon in extremely narrow circumstances where there are only two agents, a single resource, and their preferences are each other’s mirror image (i.e., if $a_{0}$ prefers $x$ over its absence, then $a_{1}$ necessarily prefers its absence over $x$.)
We now turn to the primary disproof in 4.6 where we offer an explicit counterexample to Proposition 4.5. Evidently, it would seem that the exchange mechanism prescribed by (C1) is too strong, and it has been argued that the inflexibility of a double coincidence of wants is a problem in markets where exchanges are conducted through bartering. The problem is certainly not unique to such societies and has not fully disappeared in markets where there exists a dominant medium of exchange (i.e., where it has only morphed into the need for a double coincidence of value.) To see this, we may turn to 4.6 and reinterpret $z$ as a $20\$$ bill, whereas $x$ and $y$ continue to stand for material goods. $a_{0}$ still prefers $y$ over $x$ and $x$ over $20\$$: $a_{1}$ prefers $x$ over $20\$$ and $20\$$ over $y$, while $a_{2}$ prefers $20\$$ over $y$ and $y$ over $x$. If the initial allocation were ($\big{(}\{x\},\{20\$\},\{y\}\big{)}$, then $a_{1}$ cannot purchase $x$ from $a_{0}$, who values it over $20\$$. It follows from 4.6 that no exchange occurs and the distribution is suboptimal. This is a mathematically trivial demonstration, yet its philosophical significance for economic theory lies in the observation that a market exchange should not be viewed as a private activity among only two parties. If it were, then it would be irrational for $a_{1}$ to purchase $y$ from $a_{2}$ with $20\$$, as that would make her worse off immediately. But if she chooses to do so and subsequently barters $y$ for $x$ with $a_{0}$, then everyone is made better off. Naturally, it would be more efficient if all three agents recognized this, and immediately arrived at a better distribution through a 3-party agreement. The presence of a common medium of exchange has a non-trivial effect if we assume that preferences are incomplete over resources but complete over the medium of exchange; we reject this two-sorted-ness as it is irrational and would contradict (B3). Furthermore, it is inconsistent with (B2) because the defining quality of a medium of exchange is its ability to measure the value of an item. If we can compare an item to a quantity of the medium, and compare that quantity with the quantity we derive from some other comparable item, then we should be able to compare the items themselves as though we lived in a bartering economy. [^7] Let $R$ be a set of resources such that it consists of a unique multiset of identical goods $\mathcal{N}\subseteq R$ for which every agent has quasi-linear preferences: $\forall a_{q}\in\mathbb{A}\forall A,B\in\mathcal{P}(R)\big{[}(((A\setminus(A\cap B)\subseteq\mathcal{N})\land(B\setminus(A\cap B)\subseteq\mathcal{N}))\land(|A\setminus(A\cap B)|<|B\setminus(A\cap B)|))\implies(A,B)\in W(a_{q})\big{]}.$ The existence of this numeraire does not solve the predicament since agents may freely elect to not have a price tag on particular material resources. In other words, no amount of money would be enough to compel them to sell off what they own (i.e., $\exists a_{q}\in\mathbb{A}\exists A\in\mathcal{P}(R)\forall B\subseteq\mathcal{N}[(B,A)\in W(a_{q})]$). If one were to contend that a price tag can be assigned to every bundle of resources in $\mathcal{P}(R)$ (i.e., $\forall a_{q}\in\mathbb{A}\,\forall A\in\mathcal{P}(R)\,\exists B\subseteq\mathcal{N}[(A,B)\in W(a_{q})]$) then this would call into question the value of the medium. Evidently, the numeraire’s worth is contingent on the value of the goods it can buy: it cannot be pursued as a bundle of resources in its own right, rather, there must exist some bundle of resources such that that bundle is strictly preferred over the numeraire. But this would create a vicious cycle: since there must exist a price tag for that bundle, there must also be a different corresponding bundle for that price.
Similar to a Dedekind cut, the notion of an ideal exchange (D1) is a precise solution to the predicament described in 4.6. As opposed to performing exchanges on the basis of pairwise preferences, we allow an arbitrary number of agents $(n)$ to form an $n$-party agreement iff an agent is made better off without there being another agent who is necessarily made worse off. Moreover, the distribution effectively halts once a Pareto optimal outcome is reached. From there, a proof of the Coase theorem seems almost tautological, since every ideal exchange is not merely an endeavor to reach a Pareto optimal distribution through the blueprint of agent preferences, but rather, is, in it of itself, defined in terms of Pareto optimality.
Finally, it would seem that the institutional safeguards of having well-defined ownership rights (A1) and the absence of unclaimed resources (A2) are insufficient for a proof of the Coase theorem, even if we assume that preferences are characteristically rational and transactions are frictionless. The ultimate question remains unresolved: what ought to be done given these preferences and this initial distribution? Philosophically, an ideal exchange can be characterized as a situation where every agent has sufficient entrepreneurial zeal or intellectual foresight to recognize and exploit the inefficiency of their current situation. This is especially evident for $a_{1}$ when she makes an exchange that does not reflect her authentic preferences in order to reach a more desirable, permanent outcome. It follows that the truth of the theorem rests on whether we should extend this presumption of foresight and entrepreneurial courage to any market participant at all.
Acknowledgments
The author is grateful to Noah Wang (University of St Andrews) and Professor Benedict Eastaugh (University of Warwick) for their lasting inspiration and guidance.